OpenStudy (macabbey101):

PLEASE HELP WILL FAN AND MEDAL!!! The lengths of the sides of a rectangular window have the ratio 1.2 to 1. The area of the window is 811.2 square inches. What are the dimensions of the window? a. 26 inches by 31.2 inches c. 28.48 inches by 28.48 inches b. 26 inches by 37.44 inches d. 26 inches by 52 inches

OpenStudy (macabbey101):

@hekaphage could you maybe help with this one? thank youu!

rishavraj (rishavraj):

see if breadth is x then length is 1.2x..so length * breadth = 811.2 solve for x

OpenStudy (macabbey101):

im confused how would I solve for x? @rishavraj

rishavraj (rishavraj):

so it would be \[x \times 1.2x ~=~ 811.2\]

rishavraj (rishavraj):

@macabbey101 can u solve the above equation??

OpenStudy (macabbey101):

im confused. I tried solving and the closest I got to 811 was 973. What would the equation be?

rishavraj (rishavraj):

see one side is 1.2 times the other ..so if one side is "x" the other is 1.2x... and area of rectangle is length*breadth.. given area is 811.2,.... so \[x ~\times~ 1.2x ~=~811.2\]

rishavraj (rishavraj):

so it would be \[1.2x^2 ~=~811.2\] \[x^2~=~\frac{ 811.2 }{ 1.2 }\] now proceed

OpenStudy (macabbey101):

i got 26 and -26

OpenStudy (macabbey101):

see somewhere I am going wrong with this

rishavraj (rishavraj):

so -26 is ignored coz length cant be negative ..so one side is 26...the other is supposed to be 1.2 times of 26

OpenStudy (macabbey101):

so its D?

rishavraj (rishavraj):

-_- \[26 \times 1.2 ~=~ 31.2\]

OpenStudy (macabbey101):

oh oh okay so its A im sorry I get it now I was adding 26 by 26 lol

OpenStudy (macabbey101):

Thank you!